OpenStudy (nuccioreggie):

The table shows the number of students in a class who play on a school team and/or are members of a school club: Play on a School Team Do Not Play on a School Team Total Club Member 12 ? 15 Not a Club Member 11 10 21 Total 23 13 36 What is the missing number in the table? 3 5 25 27

OpenStudy (michele_laino):

the requested number \(x\) is such that \(12+x=15\)

OpenStudy (nuccioreggie):

C

OpenStudy (michele_laino):

or: \(x+10=13\)

OpenStudy (nuccioreggie):

OH A

OpenStudy (michele_laino):

are you sure? what is \(15-12=13-10=...?\)

OpenStudy (michele_laino):

correct! It is option A